# Look Before you Leap: Estimating LLM Benchmark Scores from Descriptions

### v1

@@ Abstract
Progress in large language models is constrained by an evaluation bottleneck: build a benchmark, evaluate models and settings, then iterate. We therefore ask a simple question: can we forecast outcomes before running any experiments? We study text-only performance forecasting: estimating a model’s score from a redacted task description and intended configuration, with no access to dataset instances. To support systematic study, we curate Precog, a corpus of redacted description–performance pairs spanning diverse tasks, domains, and metrics. Experiments show the task is challenging but feasible: models equipped with a retrieval module that excludes source papers achieve moderate prediction performance with well-calibrated uncertainty, reaching mean absolute error as low as $8.7$ on the Accuracy subset at high-confidence thresholds. Our analysis indicates that stronger reasoning models engage in diverse, iterative querying, whereas current open-source models lag and often skip retrieval or gather evidence with limited diversity. We further test a zero-leakage setting—forecasting on newly released datasets or experiments before their papers are indexed—where GPT-5 with built-in web search still attains non-trivial prediction accuracy. Overall, our corpus and analyses offer an initial step toward open-ended anticipatory evaluation, supporting difficulty estimation and smarter experiment prioritization. [^1] Our data and code and are available in https://github.com/JJumSSu/PRECOG
@@ 1 Introduction
Modern LLM development is evaluation-centric: define a benchmark, sweep models and configurations, iterate (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023; Team et al., 2023). As benchmarks proliferate, exhaustive curation and full sweeps demand substantial compute, researcher time, and funding (Mazumder et al., 2023). We therefore seek an ex ante performance forecaster to help designers decide what to build and where to measure. For example, consider a team proposing a binary “solver-choice” task on an ordinary differential equations benchmark (Figure 1; Gaonkar et al., 2025). Before annotating a single example, they can ask a concrete what-if: will GPT-4.1 (Achiam et al., 2023) without CoT (Wei et al., 2022) land nearer 30 or 60 Accuracy points? A forecast like this has the potential to quickly inform scope and spend—how large a pilot to run, which baselines to include, and whether the task merits full-scale annotation.
Figure 1: Precog pairs a redacted task/dataset description with its target metric. Given the inputs (a redacted description of dataset and an experimental configuration (Gaonkar et al., 2025)), the goal is to predict the target metric. The regressor conditions on performance-relevant details and may retrieve corroborating evidence from the literature.
Two adjacent lines of work on ML model performance prediction address related goals: (i) missing-entry prediction, which infers unobserved cells from observed ones (Xia et al., 2020; Ye et al., 2021; 2023; Zhao et al., 2025); and (ii) efficient evaluation, which extrapolates full-benchmark performance from a small set of pilot runs (Maia Polo et al., 2024; Pacchiardi et al., 2024; Anugraha et al., 2025). Both assume seed measurements or closely related tasks and typically rely on engineered numeric features. In contrast, we target a zero-example, open-ended setting in which retrieval is driven by natural-language task descriptions, supporting researchers who are creating new datasets or designing new experiments. Prior work has also explored the capability of LLMs to forecast the outcomes of future geopolitical events in online prediction markets (Karger et al., 2024; Halawi et al., 2024), however to the best of our knowledge, this is the first work to explore the capability of LLMs to predict their own performance on a new task based only on a text description.
@@
To study this setting systematically, we curate Precog, a corpus of arXiv-mined experimental records (Park et al., 2025) that, for each record, pairs the dataset-source paper (introducing the benchmark) with the result-reporting paper (reporting the score). From these sources, we synthesize redacted descriptions that preserve task and setup details while masking dataset names and numeric results. At retrieval time, we exclude each record’s own source papers to limit leakage. To further probe residual leakage risks (Tian et al., 2024), we also evaluate a streaming regime that ingests arXiv releases and makes predictions before they are indexed by search engines.
We then evaluate zero-shot LLM regressors, optionally augmented with literature search. Given a redacted description, the model may iteratively issue evidence-seeking queries (ReAct-style (Yao et al., 2023)) and outputs a point estimate alongside a self-reported confidence (Yoon et al., 2025). We assess calibration by sweeping a confidence threshold and reporting the coverage–MAE trade-off, and we analyze search tool-use behavior.
On Precog, a proprietary reasoning model with literature retrieval attains MAE $14.0$ across seven metric groups and $8.7$ on the Accuracy subset at high-confidence thresholds, with reasonably calibrated uncertainty—indicating the task is hard but tractable. The proprietary reasoning model employs diverse, iterative querying, whereas current open-source models underperform and exhibit shallower tool use. In streaming prediction, GPT-5 (OpenAI, 2025) with built-in web search delivers competitive, though imperfect, performance on newly released datasets.
Altogether, our study shows that LLM-based regression models, which predict model performance from text descriptions, can help inform benchmark design, prioritize configurations, determine dataset sizes, and guide pilot experiment planning—while reducing the compute and time required for exploratory analysis. We release Precog, a redacted-description benchmark grounded in scientific literature, along with the accompanying prompts, protocols, and experimental code. We hope these resources and our evaluation methodology will support future advances in predictive model development and evaluation.
@@ 2 Related Work
Prior work predicts missing table entries from observed cells plus side information, using matrix-completion or response-surface models (Achille et al., 2019; Xia et al., 2020; Chen & Ritter, 2021; Ye et al., 2021; 2023; Rabinovich et al., 2023; Shi et al., 2025; Zhao et al., 2025). These methods typically assume (i) seed results on related tasks, (ii) stable task identities, and (iii) overlap that enables interpolation. A complementary line of work targets efficient evaluation: execute a small subset of measurements and extrapolate full-benchmark performance via learning-curve or proxy modeling (Maia Polo et al., 2024; Pacchiardi et al., 2024; Anugraha et al., 2025). These approaches still require running some evaluations on the target (or closely related) benchmark.
@@
Recent work explores LLMs as regressors in structured experimental settings, such as AutoML or black-box optimization (Song et al., 2024; Vacareanu et al., 2024; Nguyen et al., 2024; Requeima et al., 2024; Tang et al., 2024). OmniPred (Song et al., 2024), for instance, trains LLMs on millions of hyperparameter optimization trajectories from Google’s internal Vizier database and demonstrates strong regression accuracy across diverse tasks. However, these approaches rely on structured parameterizations (e.g., key–value hyperparameter strings) and assume access to seed trials or finetuning on the target task. In contrast, we study a zero-example setting where the sole input is a free-form natural-language description of a novel dataset and evaluation protocol, supporting researchers who have not yet collected any results.
@@
Prior work extracts structured result tuples (e.g., task, dataset, metric) from scientific articles (Singh et al., 2019; Hou et al., 2019; Kardas et al., 2020; Yang et al., 2022; Bai et al., 2023; Şahinuç et al., 2024; Park et al., 2025). Precog builds on these tuples: for each experimental record, we synthesize a richer natural-language description that augments the tuple with additional context.
@@ 3 Task
Given a redacted description $x_{i}$, an LLM regressor $f_{\theta}$ predicts the normalized target as $\hat{y}_{i}=f_{\theta}(x_{i},Z_{i})$, where $Z_{i}$ is an optional set of retrieved evidence. We obtain $Z_{i}$ by querying a time-bounded literature corpus $\mathcal{C}$ with a top-$k$ retriever $R_{k}$ on a query formed from $x_{i}$, i.e., $Z_{i}=R_{k}(q(x_{i}),\,\mathcal{C})$. Setting $k=0$ yields the description-only variant $f_{\theta}(x_{i},\varnothing)$. At inference, $f_{\theta}$ may iterate in a ReAct-style loop—reason (Yao et al., 2023), issue $q(\cdot)$, retrieve $Z_{i}$, refine—and then output $\hat{y}_{i}$. For example, for the open-domain QA description above, $f_{\theta}$ can retrieve prior exact-match (EM) ranges for similar setups and use them to refine $\hat{y}_{i}$.
@@
Following prior work (Srinivasan et al., 2021; Ahuja et al., 2022; Jawahar et al., 2024), we evaluate $f_{\theta}$ with two standard criteria: mean absolute error (MAE) and the Pearson correlation $\mathrm{r}$ between $y_{i}$ and $\hat{y}_{i}$. MAE (lower is better) measures absolute accuracy across heterogeneous metrics, while $\mathrm{r}$ (higher is better) assesses whether predictions preserve the relative ordering of tasks/configurations. Used together, they capture both value accuracy and trend fidelity despite differences in metric scales and variances—an evaluation protocol that is standard for performance prediction.
Task Design
Work Text-Desc. Feat. Retrieval Runs
Xia et al. (2020)
Ye et al. (2021)
Ye et al. (2023)
Maia Polo et al. (2024)
Pacchiardi et al. (2024)
Anugraha et al. (2025)
Precog (ours)
Table 1: Comparison of performance-prediction schemas. Text-Desc.: free-form descriptions. Feat.: engineered numerics features. Retrieval: literature search at prediction time. Runs: target model run on a subset.
Difference from Prior Work
Table 1 contrasts our setting with prior work: (i) it is literature-grounded and open-ended, requiring retrieval rather than seed-set evaluation; and (ii) it predicts from free-form natural-language descriptions rather than engineered numeric features. Operationally, you can onboard new tasks without running target models or maintaining seed evaluations. Free-form descriptions, augmented by retrieval from external evidence sources, capture protocol nuances and handle open-ended variation, while yielding traceable evidence for human-in-the-loop review.
@@ 4 Data Curation
To evaluate at scale, we need a large number of diverse and realistic, open-ended pairs $(x_{i},y_{i})$. Rather than synthesizing experimental configurations and running automated experiments, which is both costly and unlikely to capture the kinds of problems practitioners actually investigate, we semi-automatically curate experimental configurations and their reported results directly from the scientific literature. Our pipeline harvests redacted task descriptions and reported scores from published papers; humans intervene only for precision-critical steps. We follow the pipeline in Figure 2: (i) collect arXiv experimental records (section 4.1); (ii) define the record provenance and paper types (section 4.2); (iii) recover each record’s sources (section 4.3); (iv) extract a schema-aligned, redacted description (section 4.4); and (v) perform validation and quality control (section 4.5).
@@ 4.1 Source Corpus
We build off of LLMEvalDB (Park et al., 2025), a structured arXiv-mined repository (2023–2024) dominated by non-finetunable, proprietary models—suited to isolating dataset/task effects with less model/method variance. Each experimental record is a table cell with the dataset name, configuration context, and reported metric. Values are normalized to define $y_{i}$. We consider seven metrics with natural [0,100] scales: Accuracy, F1, Recall, Precision, ROUGE, BLEU, and EM.
@@ 4.2 Record Provenance and Paper Types
Each record is grounded in two literature sources: the result paper, which reports the score, and the dataset paper, which introduced the dataset. These may coincide or differ. For example, Paper A introduces a QA dataset and reports accuracy for several models (result paper = dataset paper). Later, Paper B evaluates a new prompting/evaluation protocol on the same dataset under a revised setup (result paper $\neq$ dataset paper) (Park et al., 2021; Rangapur & Rangapur, 2024).
@@ 4.3 Dataset Source Paper Retrieval
To compile faithful descriptions for each record, we first collect its sources. LLMEvalDB (Park et al., 2025) provides the result papers. When the result paper differs from the dataset paper, we automatically recover the latter by inferring the citation with an LLM, resolving it to an arXiv ID, downloading the PDF, and verifying the dataset name/title. We then run a rule-based check to confirm that the PDF explicitly mentions the dataset (string match). Because source fidelity is critical, any remaining errors are corrected via targeted human edits.
@@ 4.4 Extracting Redacted Input Description
Given the sources for the result and dataset papers, an extractor LLM produces a schema-aligned, redacted description $x_{i}$ that summarizes the task (problem definition and input/label format), data collection (provenance, domain, collection protocol), evaluation (metric and protocol), difficulty cues (number of classes, long context, reasoning depth), the relevant subset/split, the prompting strategy (zero/few-shot, CoT), and any other record-specific settings (context-window, tool/retrieval use), when reported; see the Appendix D for illustrative examples.
We systematically remove the dataset names, performance numbers, and statistics. To further reduce extraction-time leakage, a second LLM rewriting pass is applied and the same checks are re-enforced, yielding faithful, self-contained, anonymized inputs $x_{i}$.
Figure 2: Overview of the Precog curation pipeline. For each record, we combine the result paper (Mitchell et al., 2023), which reports the score and specifies the evaluation protocol and configuration, with the dataset paper (Moskvichev et al., 2023), which provides the task description, to produce a schema-aligned input. The score $0.30$ reported is normalized to serve as the target.
@@ 4.5 Data Statistics and Quality Control
Total instances 767
result paper == dataset paper 504
result paper $\neq$ dataset paper 263
Unique counts
Unique datasets 528
result paper $\cup$ dataset paper 631
@@
Table 2: Precog statistics and human validation of extracted descriptions.
We construct a bounded yet diverse corpus via stratified sampling over source papers, enforcing caps per paper and per dataset plus per-source quotas to prevent any single source from dominating. This yields 767 instances spanning 528 distinct datasets and 631 unique papers, covering many datasets and configurations rather than a single benchmark (Ye et al., 2023). Rather than scaling the number of instances—which would be costly for reasoning models—we focused on curating diverse and high-quality examples using the GPT-5 (OpenAI, 2025) API (\$200) and 36 hours of human validation.  [^2] We also release $\sim 9\mathrm{k}$ pre-extraction intermediate sources to support scaling evaluation and training.
Because Precog is LLM-assisted, we add fidelity safeguards: (i) humans verify and correct each record’s dataset–source link; (ii) dataset names are canonicalized and we sample over unique source pairs to promote diversity; and (iii) metrics and related fields inherited from LLMEvalDB (Park et al., 2025) exhibit near-perfect extraction accuracy. We further audited 30 random descriptions under a fixed protocol—checking anonymization (binary: no dataset names or performance numbers), schema coverage (5-point), and source support vs. the PDFs (5-point). Results (Table 2) show well-anonymized, accurate, and comprehensively grounded descriptions, suitable as schema-aligned, leakage-resistant inputs for text-only performance forecasting.

### v2

@@ Abstract
Progress in large language models is constrained by an evaluation bottleneck: build a benchmark, run models, then iterate. We ask a question: can we forecast outcomes before running any experiments to inform earlier study design? For example, a team building an AI assistant for a certain task can estimate whether expected performance is around 50 or closer to 80, evidence that supports whether to proceed to a pilot study, how to scope it, and how to allocate resources. We study text-only performance forecasting, where a model predicts a score from a redacted task description and intended configuration, with no access to dataset instances. To support systematic study, we curate Precog, a corpus of redacted description–performance pairs spanning diverse tasks, domains, and metrics. We scrape task and configuration descriptions from arXiv, yielding 2,290 instances covering 1,519 papers, and construct a leakage-free test split using papers published after the evaluated models’ knowledge cutoff. Experiments show the task is challenging but feasible: reasoning models achieve moderate prediction performance with well-calibrated uncertainty, reaching mean absolute error as low as $9.9$ at high-confidence thresholds. We further test a zero-leakage setting—forecasting on newly released datasets or experiments before their papers are indexed—where GPT-5 with built-in web search still attains non-trivial prediction accuracy. Overall, our corpus and analyses offer an initial step toward open-ended anticipatory evaluation, supporting difficulty estimation and smarter experiment prioritization. [^1] Our data and code and are available in https://github.com/JJumSSu/PRECOG
@@ 1 Introduction
Modern LLM development is evaluation-centric: define a benchmark, evaluate models and configurations, iterate (brown2020language; achiam2023gpt; touvron2023llama; team2023gemini). As benchmarks proliferate, exhaustive curation and full experiments demand substantial compute, researcher time, and funding (mazumder2023dataperf). We therefore seek an ex ante performance forecaster to help designers decide what to build and where to measure. For example, consider a team proposing a binary “solver-choice” task on an ordinary differential equations benchmark (Figure 1; gaonkar2025sciml). Before annotating a single example, they can ask a concrete what-if: will GPT-4.1 (achiam2023gpt) without CoT (wei2022chain) land nearer 30 or 60 Accuracy points? A forecast like this has the potential to quickly inform scope and spend—how large a pilot to run, which baselines to include, and whether the task merits full-scale annotation.
Figure 1: Precog pairs a redacted task/dataset description with its target metric. Given the inputs (a redacted description of dataset and an experimental configuration (gaonkar2025sciml)), the goal is to predict the target metric. The regressor conditions on performance-relevant details and may retrieve corroborating evidence from the literature.
Two adjacent lines of work on ML model performance prediction address related goals: (i) missing-entry prediction, which infers unobserved cells from observed ones (xia2020predicting; ye2021finegrained; ye2023predictable; zhaocan); and (ii) efficient evaluation, which extrapolates full-benchmark performance from a small set of pilot runs (maiapolo2024tinybenchmarks; pacchiardi2024hundred; anugraha2025proxylm). Both assume seed measurements or closely related tasks and typically rely on engineered numeric features. In contrast, we target a zero-example, open-ended setting in which retrieval is driven by natural-language task descriptions, supporting researchers who are creating new datasets or designing new experiments. Prior work has also explored the capability of LLMs to forecast the outcomes of future geopolitical events in online prediction markets (kargerforecastbench; halawi2024approaching), however to the best of our knowledge, this is the first work to explore the capability of LLMs to predict their own performance on a new task based only on a text description.
@@
To enable systematic study, we curate Precog, an arXiv-mined corpus of experimental records (park2025can) that, for each entry, links the benchmark’s source paper (introducing the dataset) to the paper reporting results on it. From these papers we extract task and configuration descriptions, yielding 2,290 instances spanning 1,519 papers, and we construct a leakage-free test split by restricting to papers published after the evaluated models’ knowledge cutoff (2025 January-September). From these sources, we synthesize redacted descriptions that preserve task and setup details while masking dataset names and numeric results. To further probe residual leakage risks (tian2024scicode), we also evaluate a streaming regime that ingests arXiv releases and makes predictions before they are indexed by search engines.
We then evaluate zero-shot LLM regressors, optionally augmented with literature search. Given a redacted description, the model may iteratively issue evidence-seeking queries (ReAct-style (yao2023react)) and outputs a point estimate alongside a self-reported confidence (yoon2025reasoning). We assess calibration by sweeping a confidence threshold and reporting the coverage–MAE trade-off, and we analyze search tool-use behavior.
On Precog, a proprietary reasoning model attains an MAE of $14.6$ across seven metric groups, and $9.9$ on a high-confidence, accuracy-filtered subset, with reasonably calibrated uncertainty. We perform both reasoning and subset ablations to identify which factors drive performance prediction and to characterize its behavior. In the streaming setting, GPT-5 (openai2025gpt5) with built-in web search delivers competitive, though imperfect, performance on newly released datasets, while maintaining good calibration.
Altogether, our study shows that LLM-based regressors that predict model performance from text descriptions can inform benchmark design, prioritize configurations, determine dataset sizes, and plan pilot experiments, while reducing compute and turnaround time for exploratory analysis. We release Precog, a redacted-description benchmark grounded in the scientific literature, together with prompts, protocols, and code. We hope these resources and our evaluation methodology will support future advances in predictive evaluation and model selection.
@@ 2 Related Work
Task Design
Work Text. Feat. Runs Idea Task Reg.
xia2020predicting
ye2021finegrained
ye2023predictable
maiapolo2024tinybenchmarks
pacchiardi2024hundred
anugraha2025proxylm
wen2025predicting
Precog (ours)
Table 1: Comparison of performance-prediction schemas. Text.: task descriptions; Feat.: numeric features; Runs: running the model on a subset of the benchmark; Idea: predicts unseen method’s performance; Task: predicts performance on an unseen task; Reg.: (Regression) predicts an metric value.
Prior works primarily fall into two categories. (1) Missing-entry prediction: infer unobserved cells from observed ones using matrix-completion or response-surface models (achille2019task2vec; xia2020predicting; chen2021model; ye2021finegrained; ye2023predictable; rabinovich2023predicting; shi2025importance; zhaocan); these assume seed results, stable task identities, and overlap for interpolation. (2) Efficient evaluation: run a small subset of trials and extrapolate full-benchmark performance (maiapolo2024tinybenchmarks; pacchiardi2024hundred; anugraha2025proxylm); these still require executing target (or closely related) runs. Finally, we note contemporaneous work by wen2025predicting, which aims to predict which of two novel research ideas is likely to perform better on a set of existing benchmarks. In contrast, our work demonstrates that LLMs can predict the absolute performance of a fixed model on a novel benchmark. See Table 1 for a detailed comparison.
@@
LLMs have been used as regressors in structured settings (AutoML/black-box optimization) (song2024omnipred; vacareanu2024words; nguyen2024predicting; requeima2024llm; tang2024understanding), often relying on key–value hyperparameter strings and access to seed trials or finetuning on the target task. In contrast, we study a zero-run, description-only regression problem where the sole input is a free-form specification of a novel dataset and evaluation protocol. Our task involves forecasting, specifically, predicting the outcomes of ideas before conducting real experiments. (zou2022forecasting; schoenegger2023large; kargerforecastbench; abolghasemi2025humans; moussa2025scholareval). Inspired by these works, our work demonstrates the potential of reasoning LLMs in forecasting the outcomes of evaluation results on unseen datasets.
@@
Prior work extracts structured result tuples (e.g., task, dataset, metric) from scientific articles (singh2019automated; hou2019identification; kardas2020axcell; yang2022telin; bai2023schema; csahinucc2024efficient; park2025can). Precog builds on these tuples: for each experimental record, we synthesize a richer natural-language description that augments the tuple with additional context.
@@ 3 Task
Given a redacted description $x_{i}$, an LLM regressor $f_{\theta}$ predicts the normalized target as $\hat{y}_{i}=f_{\theta}(x_{i},Z_{i})$, where $Z_{i}$ is an optional set of retrieved evidence. We obtain $Z_{i}$ by querying a time-bounded literature corpus $\mathcal{C}$ with a top-$k$ retriever $R_{k}$ on a query formed from $x_{i}$, i.e., $Z_{i}=R_{k}(q(x_{i}),\,\mathcal{C})$. Setting $k=0$ yields the description-only variant $f_{\theta}(x_{i},\varnothing)$. At inference, $f_{\theta}$ may iterate in a ReAct-style loop—reason (yao2023react), issue $q(\cdot)$, retrieve $Z_{i}$, refine—and then output $\hat{y}_{i}$. For example, for the open-domain QA description above, $f_{\theta}$ can retrieve prior exact-match (EM) ranges for similar setups and use them to refine $\hat{y}_{i}$.
@@
Following prior work (srinivasan2021predicting; ahuja2022multi; jawahar2024llmpp), we evaluate $f_{\theta}$ with two standard criteria: mean absolute error (MAE) and the Pearson correlation $\mathrm{r}$ between $y_{i}$ and $\hat{y}_{i}$. MAE (lower is better) measures absolute accuracy across heterogeneous metrics, while $\mathrm{r}$ (higher is better) assesses whether predictions preserve the relative ordering of tasks/configurations. Used together, they capture both value accuracy and trend fidelity despite differences in metric scales and variances—an evaluation protocol that is standard for performance prediction.
@@ 4 Data Curation
To evaluate at scale, we need a large number of diverse and realistic, open-ended pairs $(x_{i},y_{i})$. Rather than synthesizing experimental configurations and running automated experiments, which is both costly and unlikely to capture the kinds of problems practitioners actually investigate, we automatically curate experimental configurations and their reported results directly from the scientific literature. Our pipeline harvests redacted task descriptions and reported scores from published papers. We follow the pipeline in Figure 2: (i) collect arXiv experimental records (section 4.1); (ii) define the record provenance and paper types (section 4.2); (iii) recover each record’s sources (section 4.3); (iv) extract a schema-aligned, redacted description (section 4.4); and (v) perform validation and quality control (section 4.5).
@@ 4.1 Source Corpus
We build on LLMEvalDB (park2025can), focusing on its 2023–2024 entries and 2025 entries from January to September, which are dominated by non-finetunable, proprietary models. This choice reduces variance from model/method tuning and better isolates dataset/task effects. Moreover, because current LLMs generally have a 2024 knowledge cutoff, restricting the 2025 split to post-cutoff papers mitigates leakage: the reported metrics could not have appeared in the training data. Each experimental record is a table cell with the dataset name, configuration context, and reported metric. Values are normalized to define $y_{i}$. We consider seven metrics with natural [0,100] scales: Accuracy, F1, Recall, Precision, ROUGE, BLEU, and EM.
@@ 4.2 Record Provenance and Paper Types
Each record is grounded in two literature sources: the result paper, which reports the score, and the dataset paper, which introduced the dataset. These may coincide or differ. For example, Paper A introduces a QA dataset and reports accuracy for several models (result paper = dataset paper). Later, Paper B evaluates a new prompting/evaluation protocol on the same dataset under a revised setup (result paper $\neq$ dataset paper) (park2021faviq; rangapur2024battle).
@@ 4.3 Dataset Source Paper Retrieval
To compile faithful descriptions for each record, we first collect its sources. LLMEvalDB (park2025can) provides the result papers. When the result paper differs from the dataset paper, we automatically recover the latter by inferring the citation with an LLM, resolving it to an arXiv ID, downloading the PDF, and verifying the dataset name/title. We then run a rule-based check to confirm that the PDF explicitly mentions the dataset (string match).
@@ 4.4 Extracting Redacted Input Description
Given the sources for the result and dataset papers, an extractor LLM produces a schema-aligned, redacted description $x_{i}$ that summarizes the task (problem definition and input/label format), data collection (provenance, domain, collection protocol), evaluation (metric and protocol), difficulty cues (number of classes, long context, reasoning depth), the relevant subset/split, the prompting strategy (zero/few-shot, CoT), and any other record-specific settings (context-window, tool/retrieval use), when reported; see the Appendix E for illustrative examples.
When the retrieved sources do not actually cover the experimental record (e.g., due to retrieval errors from 4.3), we discard the case by having the extractor return failed, thereby correcting residual retrieval mistakes. For the remaining records, we remove dataset names, numerical results, and summary statistics to produce faithful, self-contained, anonymized.
Figure 2: Overview of the Precog curation pipeline. For each record, we combine the result paper (mitchell2023comparing), which reports the score and specifies the evaluation protocol and configuration, with the dataset paper (moskvichev2023conceptarc), which provides the task description, to produce a schema-aligned input. The score $0.30$ reported is normalized to serve as the target.
@@ 4.5 Data Statistics and Quality Control
total instances 767 1,523
result paper == dataset paper 504 549
result paper $\neq$ dataset paper 263 974
Paper counts
unique result paper 443 579
unique dataset paper 471 616
result paper $\cup$ dataset paper 631 956
@@
Table 2: Precog statistics by collection period. Validation was conducted on 2023–2024 period.
We construct a bounded yet diverse corpus via stratified sampling over source papers, enforcing caps per paper and per dataset plus per-source quotas to prevent any single source from dominating. This yields 2,290 instances of records spanning distinct 1,019 dataset papers and 1,022 unique result papers, covering many datasets and configurations rather than a single benchmark (ye2023predictable). Because Precog is LLM-assisted, we further audited 30 random descriptions under a fixed protocol—checking anonymization (binary: no dataset names or performance numbers), schema coverage (5-point), and source support vs. the PDFs (5-point). Results (Table 2) show well-anonymized, accurate, and comprehensively grounded descriptions, suitable as schema-aligned, leakage-resistant inputs for text-only performance forecasting.